9) What is the gram solubility of CaF2 in 100.0 mL of pure water at 25 degree C. The solubility product constant of CaF2 is 3.9*10^-11. The molar mass is 78.1 g/mol.
Answer #1

Caf2 ; - ksp = [(alt] (6-72 if solubility in Central) is s Ksp = sx (25)? S= (3-9410)= 2.136x1024M in 100 ml mass of solubl